A sample of 262 adults is selected. The adults are classified according to their voter registration status and their preferred source of current events information. The results are given in the contingency table below.

What is the relative frequency of adults in the sample who are registered to vote and prefer to obtain their information through television?

Round your answer to two decimal places.

                                                                               Preferred Source of Information

                                            Television Newspapers Radio   Internet Sites

Registered

36

36

27

25

Non - Registered

18

47

30

43
Answer #1

total frequency = 36+36+27+25+18+47+30+43 = 262

frequency for adults in the sample who are registered to vote and prefer to obtain their information through television = 36

Relative frequency = (adults in the sample who are registered to vote and prefer to obtain their information through television)/(total frequency)

Relative frequency = 36/262 = 0.14